Run Meeting: August 12th, 2020
Summary of past 24h:
Tuesday, August 11 to Wednesday, August 12
Beam maintenance and restoration during day and swing shift. Magnet control software updated.
5-pass data taking at 200 nA on deuterium (about 20 M events).
Deuterium target purged opportunistically at 8 PM on Tuesday.
GEM amplification HV is back at 388 to 390 V.
New trigger configuration for FT test in use.
Plan for next 24 hours
Data taking with 200 nA beam.
Change D2 gas bottle.